# Augmenting the action space with conventions to improve multi-agent cooperation in Hanabi

### v2

@@ Abstract
The card game Hanabi is considered a strong medium for the testing and development of multi-agent reinforcement learning (MARL) algorithms, due to its cooperative nature, hidden information, limited communication and remarkable complexity. Previous research efforts have explored the capabilities of MARL algorithms within Hanabi, focusing largely on advanced architecture design and algorithmic manipulations to achieve state-of-the-art performance for various number of cooperators. However, this often leads to complex solution strategies with high computational cost and requiring large amounts of training data. For humans to solve the Hanabi game effectively, they require the use of conventions, which often allows for a means to implicitly convey ideas or knowledge based on a predefined, and mutually agreed upon, set of “rules” or principles. Multi-agent problems containing partial observability, especially when limited communication is present, can benefit greatly from the use of implicit knowledge sharing. In this paper, we propose a novel approach to augmenting an agent’s action space using conventions, which act as a sequence of special cooperative actions that span over and include multiple time steps and multiple agents, requiring agents to actively opt in for it to reach fruition. These conventions are based on existing human conventions, and result in a significant improvement on the performance of existing techniques for self-play and cross-play across various number of cooperators within Hanabi.
@@
Multi-agent Reinforcement Learning, Deep Learning, Hanabi, Agent Cooperation, Cross-play
@@ 1 Introduction
When another agent controlling a different vehicle is introduced into the environment, the need for communication to overcome the problem of partial observability becomes apparent, since this agent will have a different viewpoint of the environment. These viewpoints will have some overlap, but will also contain important information that the agents will have to share, such as obscured objects, accidents, obstacles, pedestrians, etc., in order to safely navigate the road. Even when the agents are able to communicate, it is infeasible to share and process all the high quality information between agents, especially in scenarios where quick reaction times are crucial. This communication problem becomes even more difficult if we consider situations where more than two vehicles must coordinate effectively, for example a large intersection, road construction, an accident that lead to congestion, or a busy city.
@@
One of the most straightforward solutions to multi-agent reinforcement learning (MARL) is the combination of deep Q-networks (DQNs) and independent Q-learning [6], where each agent independently and simultaneously learns its own action-value function (or Q-values [^1] Action values ($Q$) refer to the value assigned to a certain action $a$ within a given state $s$.) while interacting with the same environment. However, this strategy does not fair well when paired with partial observability [7, 8]. Hausknecht and Stone [9] have shown that recurrent neural networks offer an improved solution to MARL problems containing partial observability. These networks incorporate a built-in short term memory over which experiences are unrolled (or combined) to form longer sequences. This is often combined with deep Q-learning’s feed forward neural network to produce deep recurrent Q-learning [9].
@@
Various research efforts follow a similar path as Hausknecht and Stone, focusing on architectural advancements through the means of neural network layer manipulations, and complex algorithms to improve the estimation and assignment of action-values or policies [10, 11, 12]. This often leads to convoluted solution strategies that are difficult to implement and computationally expensive [13, 14]. An alternative solution strategy would be to reconsider the problem dynamics and discover ways of incorporating existing domain knowledge into RL algorithms. Existing domain knowledge can be incorporated into RL and MARL using reward shaping [15, 16, 17], which focuses on manipulating the reward signal to encourage certain behaviour. Alternatively, agents can use state-space augmentations, such as auxiliary tasks [18], to imply or extract additional information (or features) of the problem setting. Sutton et al. [19] have shown that options offer an additional strategy for incorporating domain knowledge into RL by changing the action space of an agent. Options allow an agent to solve problems on a higher level by extending the action space to include advanced temporal actions that range over multiple time steps (also referred to as macro-actions[19].
@@
Games are often used as a medium for testing and evaluating RL algorithms, since they incorporate real-world problems with well-defined rules and a clear metric for measuring performance, where examples include Go [20], Backgammon [21], and Dota 2 [22]. Hanabi is a cooperative card game containing partial observability and limited communication, requiring players to logically reason over the intentions and actions of their cooperators, a concept known as theory of mind [23]. Human players require the development of conventions [^2] Philosopher David Lewis defines conventions as an arbitrary, self-perpetuating solution to a recurring coordination problem [24]. in order to solve the problem effectively. Conventions have evolved with humanity throughout the years and range from driving on a certain side of the road to social conventions or norms [24].
@@
Conventions allow for additional implicit communication through actions and mutually agreed upon “rules”, thereby overcoming the communication restrictions of the communication channel. Similar to the Hanabi problem, when humans tackle the problem of controlling vehicles on roads, conventions are common practice and ensure effective cooperation. Examples include: driving on a certain side of the road [^3] Philosopher David Lewis discusses this topic in depth, and places particular focus on how this affects other aspects of life, such as people tending to keep to the same side as their road conventions when walking in a busy street [24].; keep left, pass right (or vice versa); who has rite of passage at a T-intersection or four-way intersection (and how this changes based on the presence of pedestrians); flashing of headlights while stationary to indicate giving right of way/yielding; flashing of headlights while moving to indicate danger ahead, or be wary of potential danger ahead; flashing of hazards to indicate an emergency; a quick flash of hazards to say thank you; and many more. Furthermore, this highlights the importance for artificial agents to learn existing human conventions, since there will exist a period of time when autonomous agents must navigate these complex scenarios alongside humans, and the agents will have to be able to convey and receive information effectively without using their agent-specific communication channel.
@@
Most research efforts on the Hanabi problem use complex architecture design with advanced algorithms, often focusing on estimating and calculating the beliefs of other agents [25, 13]. However, one avenue of the Hanabi problem yet to be explored is that of conventions, and how to incorporate existing human conventions into MARL solution strategies. In this paper, we propose a method to incorporate human conventions into MARL through the use of artificial conventions, which acts as cooperative actions that span over and include multiple time steps and multiple agents, and show how it significantly improves on the performance of MARL agents within Hanabi for self-play [26] as well as cross-play [27] scenarios. Our approach shares similarities with options, due to the multi-time step extension of actions, however is fundamentally built on a different concept.
@@ 1.1 Related Work
Hanabi was first proposed as a viable medium and frontier for MARL by Bard et al. [36], with focus placed on the philosophical ideas and challenges found within the problem setting and how they translate to real-world scenarios. Bard et al. conducted initial tests for three different MARL algorithms, namely MA-Rainbow, actor-critic-Hanabi-agent (ACHA) [36], and Bayesian action decoder (BAD) [25], and compared them to state-of-the-art handcrafted bots, such as HatBot [37] and WTFWThat [38]. In the most difficult scenario of five players, the MARL solutions achieved an acceptable score of 16.8/25, however this still significantly falls behind the best handcrafted bot able to achieve a score of 24.89/25.
@@
The simplified action decoder (SAD), introduced by Hu and Foerster [13], aims to solve a similar goal as the BAD algorithm, but with the main benefit of reducing the effect exploration actions have on the public beliefs of agents [13]. They improve on the performance of the BAD algorithm, while also extending these concepts to more than just two players. Recurrent neural networks [9] are used to track the public beliefs and an auxiliary task [18] is trained to predict if a card is playable, discardable or unknown. Hu and Foerster found that these auxiliary tasks only benefited the two player scenario while drastically hurting the 3–5 player performance [13]. They are able to achieve good performance for 2–5 players in Hanabi, and close the gap between RL agents and handcrafted solutions by achieving a score of 22.06/25 in five player Hanabi. However, the SAD algorithm is computationally expensive, requiring a minimum of 40 CPU cores, 2 high-end GPUs and 256 gigabytes of RAM, while still requiring billions of samples and a wall time of 72 hours [13].
@@
Lerer et al. [14] further builds on the concept of the SAD algorithm by introducing multi-agent tree search (MATS) [39], and manage to achieve state-of-the-art performance for 2–5 player Hanabi. The agents manage to beat the current best handcrafted bots in the two player scenario, with the handcrafted bots only outperforming the agents in five player by 4%. However, just like its predecessor, the SAD-MATS algorithm is computationally expensive and requires a substantial amount of training data and wall time to achieve these results [14].
@@
Hu et al. [27] also builds on the concept of the SAD algorithm, but rather than achieving state-of-the-art performance for self-play agents, they focus on the agent’s ability to cooperate with a variety of partners–a concept known as cross-play. This is achieved using other-play, an architecture for multi-agent systems specifically designed to break the ties formed between self-play agents, allowing them to produce more robust strategies. This allows the agents to cooperate with agents from different training runs (or regimes), and even human cooperators [27]. Even though the other-play agents are able to achieve effective cross-play performance when paired with a variety of agents in a two player scenario, the research does not focus on higher player counts where more advanced cooperative strategies are required. Additionally, other-play builds on the existing architecture of the SAD algorithm leading to increased computational complexity and cost [27].
@@ 1.2 Summary of Contributions
Rather than focusing on complex architecture design, this research shifts focus to incorporating existing domain knowledge of the Hanabi problem into MARL algorithms. Hanabi has an active and dedicated community with players constantly searching for new conventions and ways to reliably beat the game. These conventions are often based on principles [^4] A convention principle is a strategy, or user defined “rule”, external to the existing game rules which governs a player’s behaviour and results in a desired outcome., and have generally been standardised within the Hanabi community. Our main contribution is to show how these existing human conventions can be implemented in a MARL scenario using artificial conventions, which act as a special form of cooperative actions and can be incorporated into existing MARL algorithms using action space augmentation.
@@
Our approach shares similarities with that of multi-agent options, but addresses a problem where agents must often wait for other agents to complete their options before they can initiate a new option (which can include a cooperative option) [46]. Fundamentally, our approach is built on a different concept, due to the cooperative requirement of conventions to reach fruition, and their ability to convey ideas or intentions through implicit communication based on mutually agreed upon “rules”. Therefore, it is important to note that our approach does not introduce additional explicit communication, or additional communication channels for that matter, but rather allows for implicit communication through the nature of conventions. Furthermore, each agent forming part of the convention actively participates in every step of the convention, rather than an external policy taking control and executing over multiple time steps until the option terminates.
Ultimately, we believe our work will act as the foundation for convention discovery, i.e. the ability for agents to define their conventions as they train. Similar to how options paved the way for option discovery by showing the importance and benefits of options, conventions offer a similar argument for multi-agent cooperation in partially observable environments. Option discovery has been proven to be a difficult problem to solve in multi-agent systems [46, 47], and we believe convention discovery would offer a similar challenge. Therefore, it is important to first prove the capabilities of conventions, their implications, implementation, and limitations, especially at the hands of a difficult problem such as Hanabi.
Conventions result in a large performance uplift for existing MARL techniques, specifically multi-agent Rainbow, by significantly decreasing the training time and, in the case of 3–5 player Hanabi, increasing the converged performance. Additionally, since the agents are learning from a communal and well-defined list of principles, the agents are not just able to cooperate in self-play, but also in a cross-play setting with agents from different training runs (or regimes). Cross-play has been shown to be an important research area for MARL since it opens the door to agents being able to cooperate with never-before-seen teammates, whether they are agents from other regimes, agents with different architectures, or humans (such as autonomous agents navigating the road alongside humans) [48, 49].
@@ 2 Multi-agent Reinforcement Learning and Options
where $S$ denotes the global state space and $A$ represents the joint-action space of the $n$ agents at time $t$. The observation space $O$ consists of the local observation of each agent $i$ at time $t$ within the global state $S$ ($O^{i}_{t}=O(S;i;t)$). After $A$ is sent to the environment, the global state transitions from $S$ to $S^{\prime}$ given the state transition probability $T$, and similarly $P$ represents the observation transition probability to transition from $O$ to $O^{\prime}$. This results in the environment producing $R$, which denotes the immediate reward for each agent action that contributed to a global state change at $t$, with $\gamma$ denoting the discount factor for future reward. These rewards are either combined with equal weighting (summed), or can have their own discount factor $\gamma_{r}$, and is referred to as the forward accumulated reward [51].
In value-based RL we can use (1) to calculate the value functions ($Q$) which acts as a quantitative measure for the desirability of a state. In tabular methods these values are represented using a lookup-table, while deep RL methods use powerful function approximators to estimate these values.
@@ 2.2 Deep Q-learning
Deep Q-learning usually incorporates a policy network and a target network [53]. The policy network is used to select actions and utilises the update step, while the target network serves as a baseline when calculating the TD error. The target network is updated periodically with the policy network to reduce correlation with the target. The weights of the policy network are updated using backpropagation with the goal of minimizing the TD error, which is defined as
@@
where $\theta_{j}$ and $\theta_{j}^{-}$ are the weights of the policy-and target network, respectively, at iteration $j$ [53]. In practice the backpropagation and calculation of the TD error is usually handled by an optimizer, such as the Adam optimizer [55]. Similar to tabular Q-learning, deep Q-learning can be extended to MARL using independent deep Q-learning [6], and can also make use of a shared policy.

### v3

@@ Abstract
The card game Hanabi is considered a strong medium for the testing and development of multi-agent reinforcement learning (MARL) algorithms, due to its cooperative nature, partial observability, limited communication and remarkable complexity. Previous research efforts have explored the capabilities of MARL algorithms within Hanabi, focusing largely on advanced architecture design and algorithmic manipulations to achieve state-of-the-art performance for various number of cooperators. However, this often leads to complex solution strategies with high computational cost and requiring large amounts of training data. For humans to solve the Hanabi game effectively, they require the use of conventions, which often allows for a means to implicitly convey ideas or knowledge based on a predefined, and mutually agreed upon, set of “rules” or principles. Multi-agent problems containing partial observability, especially when limited communication is present, can benefit greatly from the use of implicit knowledge sharing. In this paper, we propose a novel approach to augmenting an agent’s action space using conventions, which act as a sequence of special cooperative actions that span over and include multiple time steps and multiple agents, requiring agents to actively opt in for it to reach fruition. These conventions are based on existing human conventions, and result in a significant improvement on the performance of existing techniques for self-play and cross-play for various number of cooperators within Hanabi.
@@
Multi-agent Reinforcement Learning, Hanabi, Agent Cooperation, Self-play, Cross-play
@@ 1 Introduction
When another agent controlling a different vehicle is introduced into the environment, the need for communication to overcome the problem of partial observability becomes apparent, since this agent will have a different viewpoint of the environment. These viewpoints will have some overlap, but will also contain important information that the agents will have to share, such as obscured objects, accidents, obstacles, pedestrians, etc., in order to safely navigate the road. Even when the agents are able to communicate, it is infeasible to share and process all the high-quality information between agents, especially in scenarios where quick reaction times are crucial. This communication problem becomes even more difficult if we consider situations where more than two vehicles must coordinate effectively, for example a large intersection, road construction, an accident that lead to congestion, or a busy city.
@@
One of the most straightforward solutions to multi-agent reinforcement learning (MARL) is the combination of deep Q-networks (DQNs) and independent Q-learning [6], where each agent independently and simultaneously learns its own action-value function (or Q-values [^1] Action values ($Q$) refer to the value assigned to a certain action $a$ within a given state $s$.) while interacting with the same environment. However, this strategy does not fair well when paired with partial observability [7, 8]. Hausknecht and Stone [9] have shown that recurrent neural networks offer an improved solution to MARL problems containing partial observability. These networks incorporate a built-in short-term memory over which experiences are unrolled (or combined) to form longer sequences. This is often combined with deep Q-learning’s feed forward neural network to produce deep recurrent Q-learning [9].
@@
Various research efforts follow a similar path as Hausknecht and Stone, focusing on architectural advancements through the means of neural network layer manipulations, and complex algorithms to improve the estimation and assignment of action-values or policies [10, 11, 12]. This often leads to convoluted solution strategies that are difficult to implement and computationally expensive [13, 14]. An alternative solution strategy would be to reconsider the problem dynamics and discover ways of incorporating existing domain knowledge into RL algorithms. Existing domain knowledge can be incorporated into RL and MARL using reward shaping [15, 16, 17], which focuses on manipulating the reward signal to encourage certain behaviour. Alternatively, agents can use state-space augmentations, such as auxiliary tasks [18, 19], to imply or extract additional information (or features) of the problem setting. Sutton et al. [20] have shown that options offer an additional strategy for incorporating domain knowledge into RL by changing the action space of an agent. Options allow an agent to solve problems on a higher level by extending the action space to include advanced temporal actions that range over multiple time steps (also referred to as macro-actions).
@@
Games are often used as a medium for testing and evaluating RL algorithms, since they incorporate real-world problems with well-defined rules and a clear metric for measuring performance, where examples include Go [21], Backgammon [22], and Dota 2 [23]. Hanabi is a cooperative card game containing partial observability and limited communication, requiring players to logically reason over the intentions and actions of their cooperators, a concept known as theory of mind [24]. Human players require the development of conventions [^2] Philosopher David Lewis defines conventions as an arbitrary, self-perpetuating solution to a recurring coordination problem [25]. in order to solve the problem effectively. Conventions have evolved with humanity throughout the ages and range from driving on a certain side of the road to social conventions or norms [25].
@@
Conventions allow for additional implicit communication through actions and mutually agreed upon “rules”, thereby overcoming the communication restrictions of the communication channel. Similar to the Hanabi problem, when humans tackle the problem of controlling vehicles on roads, conventions are common practice and ensure effective cooperation. Examples include: driving on a certain side of the road [^3] Philosopher David Lewis discusses this topic in depth, and places particular focus on how this affects other aspects of life, such as people tending to keep to the same side as their road conventions when walking in a busy street [25].; keep left, pass right (or vice versa); who has right of passage at a T-intersection or four-way intersection (and how this changes based on the presence of pedestrians); flashing of headlights while stationary to indicate giving right of way/yielding; flashing of headlights while moving to indicate danger ahead, or be wary of potential danger ahead; flashing of hazards to indicate an emergency; a quick flash of hazards to say thank you; and many more. Furthermore, this highlights the importance of artificial agents learning existing human conventions, since there will come a time when autonomous agents must navigate these complex scenarios alongside humans, and the agents will have to be able to convey and receive information effectively without using their agent-specific communication channel.
@@
Most research efforts on the Hanabi problem use complex architecture design with advanced algorithms, often focusing on estimating and calculating the beliefs of other agents [26, 13]. However, one avenue of the Hanabi problem yet to be explored is that of conventions, and how to incorporate existing human conventions into MARL solution strategies. In this paper, we propose a method to incorporate human conventions into MARL through the use of artificial conventions, which act as cooperative actions that span over and include multiple time steps and multiple agents, as well as show how it significantly improves on the performance of MARL agents within Hanabi for self-play [27] as well as cross-play [28] scenarios. Our approach shares similarities with options, due to the multi-time step extension of actions, however is fundamentally built on a different concept.
@@ 1.1 Related Work
Hanabi was first proposed as a viable medium and frontier for MARL by Bard et al. [37], with focus placed on the philosophical ideas and challenges found within the problem setting and how they translate to real-world scenarios. Bard et al. conducted initial tests for three different MARL algorithms, namely MA-Rainbow, actor-critic-Hanabi-agent (ACHA) [37], and Bayesian action decoder (BAD) [26], comparing them to state-of-the-art handcrafted bots, such as HatBot [38] and WTFWThat [39]. In the most difficult scenario of five players, the MARL solutions achieved an acceptable score of 16.8/25, however this still significantly falls behind the best handcrafted bot able to achieve a score of 24.89/25.
@@
The simplified action decoder (SAD), introduced by Hu and Foerster [13], aims to solve a similar goal as the BAD algorithm, but with the main benefit of reducing the effect exploration actions have on the public beliefs of agents [13]. They improve on the performance of the BAD algorithm, while also extending these concepts to more than just two players. Recurrent neural networks [9] are used to track the public beliefs and an auxiliary task [18] is trained to predict if a card is playable, discardable or unknown. Hu and Foerster found that these auxiliary tasks only benefited the two-player scenario while drastically hurting the 3–5 player performance [13]. They are able to achieve good performance for 2–5 players in Hanabi, and close the gap between RL agents and handcrafted solutions by achieving a score of 22.06/25 in five-player Hanabi. However, the SAD algorithm is computationally expensive, requiring a minimum of 40 CPU cores, 2 high-end GPUs and 256 gigabytes of RAM, while still requiring billions of samples and a wall time of 72 hours [13].
@@
Lerer et al. [14] further builds on the concept of the SAD algorithm by introducing multi-agent tree search (MATS) [40], and manage to achieve state-of-the-art performance for 2–5 player Hanabi. The agents manage to beat the current best handcrafted bots in the two-player scenario, with the handcrafted bots only outperforming the agents in five-player by 4%. However, just like its predecessor, the SAD-MATS algorithm is computationally expensive and requires a substantial amount of training data and wall time to achieve these results [14].
@@
Hu et al. [28] also builds on the concept of the SAD algorithm, but rather than achieving state-of-the-art performance for self-play agents, they focus on the agent’s ability to cooperate with a variety of partners–a concept known as cross-play. This is achieved using other-play, an architecture for multi-agent systems specifically designed to break the ties formed between self-play agents, allowing them to produce more robust strategies. This allows the agents to cooperate with agents from different training runs (or regimes), and even human cooperators [28]. Even though the other-play agents are able to achieve effective cross-play performance when paired with a variety of agents in a two-player scenario, the research does not focus on higher player counts where more advanced cooperative strategies are required. Additionally, other-play builds on the existing architecture of the SAD algorithm leading to increased computational complexity and cost [28].
@@ 1.2 Summary of Contributions
Rather than focusing on complex architecture design, this research shifts focus to incorporating existing domain knowledge of the Hanabi problem into MARL algorithms. Hanabi has an active and dedicated community of players constantly searching for new conventions and ways to reliably beat the game. These conventions are often based on principles [^4] A convention principle is a strategy, or user defined “rule”, external to the existing game rules which governs a player’s behaviour and results in a desired outcome., and have generally been standardised within the Hanabi player-base. Our main contribution is to show how these existing human conventions can be implemented in a MARL scenario using artificial conventions, which act as a special form of cooperative actions and can be incorporated into existing MARL algorithms using action space augmentation.
@@
Our approach shares similarities with that of multi-agent options, but addresses a problem where agents must often wait for other agents to complete their options before they can initiate a new option (which can include a cooperative option) [47]. Fundamentally, our approach is built on a different concept, due to the cooperative requirement of conventions to reach fruition, and their ability to convey ideas or intentions through implicit communication based on mutually agreed upon “rules”. Therefore, it is important to note that our approach does not introduce additional explicit communication, or additional communication channels for that matter, but rather allows for implicit communication through the nature of conventions. Furthermore, each agent forming part of a convention must actively choose to participate in every step of the convention, rather than an external policy taking control and executing over multiple time steps until the option terminates.
Ultimately, we believe our work will act as the foundation for convention discovery, i.e. the ability for agents to define their conventions as they train. Similar to how options paved the way for option discovery by showing the importance and benefits of options, conventions offer a similar argument for multi-agent cooperation in partially observable environments. Option discovery has been proven to be a difficult problem to solve in multi-agent systems [47, 48], and we believe convention discovery would offer a similar challenge. Therefore, it is important to first prove the capabilities of conventions, their implications, implementation, and limitations, especially at the hands of a difficult problem setting such as Hanabi.
Conventions result in a large performance uplift for existing MARL techniques, specifically multi-agent Rainbow, by significantly decreasing the training time and, in the case of 3–5 player Hanabi, increasing the converged performance. Additionally, since the agents are learning from a communal and well-defined list of principles, the agents are not just able to cooperate in self-play, but also in a cross-play setting with agents from different training runs (or regimes). Cross-play has been shown to be an important research area for MARL, since it opens the door to agents being able to cooperate with never-before-seen teammates, whether they are agents from other regimes, agents with different architectures, or humans (such as autonomous agents navigating the road alongside humans) [49, 50].
@@ 2 Multi-agent Reinforcement Learning and Options
where $S$ denotes the global state space and $A$ represents the joint-action space of the $n$ agents at time $t$. The observation space $O$ consists of the local observation of each agent $i$ at time $t$ within the global state $S$ ($O^{i}_{t}=\mathcal{O}(S;i;t)$). After $A$ is sent to the environment, the global state transitions from $S$ to $S^{\prime}$ given the state transition probability $T$, and similarly $P$ represents the observation transition probability to transition from $O$ to $O^{\prime}$. This results in the environment producing $R$, which denotes the immediate reward for each agent action that contributed to a global state change at $t$, with $\gamma$ denoting the discount factor for future reward. These rewards are either combined with equal weighting (summed), or can have their own discount factor $\gamma_{r}$, and is referred to as the forward accumulated reward [52].
In value-based RL we can use equation (1) to calculate the value functions ($Q$) which acts as a quantitative measure for the desirability of a state. In tabular methods these values are represented using a lookup-table, while deep RL methods use powerful function approximators to estimate these values.
@@ 2.2 Deep Q-learning
Deep Q-learning usually incorporates a policy network and a target network [54]. The policy network is used to select actions and utilises the update step, while the target network serves as a baseline when calculating the TD-error. The target network is updated periodically with the policy network to reduce correlation with the target. The weights of the policy network are updated using backpropagation with the goal of minimizing the TD-error, which is defined as
@@
where $\theta_{j}$ and $\theta_{j}^{-}$ are the weights of the policy-and target network, respectively, at iteration $j$ [54]. In practice the backpropagation and calculation of the TD-error is usually handled by an optimizer, such as the Adam optimizer [56]. Similar to tabular Q-learning, deep Q-learning can be extended to MARL using independent deep Q-learning [6], and can also make use of a shared policy.